OpenStudy (anonymous):

From a point on the ground 144 ft. from the base of a building, the angle of elevation to the top of the building is 23.5°. How tall is the building?

OpenStudy (amistre64):

how level is the ground?

OpenStudy (amistre64):



OpenStudy (anonymous):

it's trig so I am guessing a level plane which would mean it has a 90° angle at C.

OpenStudy (amistre64):

then tangent helps

OpenStudy (amistre64):

tan(t) = h/d d tan(t) = h

OpenStudy (anonymous):

Make a triangle out of the problem. Your horizontal leg is 144ft. Your vertical leg will be x because that is what you want to find. Your angle will be in between your horizontal leg and hypotenuse. Now it is time to pick a trig function that will answer your question... Tangent is opp/adj so it will incorporate what you need. Your equation is now tan23.5 = x/144. Now solve for x. .4348 = x/144 x is approximately 62.6